HETAV PATEL

Grade 11,

A machine gun can fire bullets of 50 gram at a speed of 2000m/sec the man holding the gun can exert an average force of 200 newtons against the gun .calculate the maximum number of bullets which he can fire per minute

A machine gun can fire bullets of 50 gram at a speed of 2000m/sec the man holding the gun can exert an average force of 200 newtons against the gun .calculate the maximum number of bullets which he can fire per minute

Grade:11

1 Answers

Vikas TU
7 years ago
Depends on the impulse he made:
Thus,
P = 50*2000/1000 = > 100*1 Kgm/s = > Impulse,
Impulse he made=> 200*60 =>12,000
Hence number of bullets = > 12000/100 = > 120